### v1

# Mini-Monkey: Alleviate the Sawtooth Effect by Multi-Scale Adaptive Cropping

Abstract
Recently, there has been significant interest in enhancing the capability of multimodal large language models (MLLMs) to process high-resolution images. Most existing methods focus on adopting a cropping strategy to improve the ability of multimodal large language models to understand image details. However, this cropping operation inevitably causes the segmentation of objects and connected areas, which impairs the MLLM’s ability to recognize small or irregularly shaped objects or text. This issue is particularly evident in lightweight MLLMs. Addressing this issue, we propose Mini-Monkey, a lightweight MLLM that incorporates a plug-and-play method called multi-scale adaptive crop strategy (MSAC). Mini-Monkey adaptively generates multi-scale representations, allowing it to select non-segmented objects from various scales. To mitigate the computational overhead introduced by MSAC, we propose a Scale Compression Mechanism (SCM), which effectively compresses image tokens. Mini-Monkey achieves state-of-the-art performance among 2B-parameter MLLMs. It not only demonstrates leading performance on a variety of general multimodal understanding tasks but also shows consistent improvements in document understanding capabilities. On the OCRBench, Mini-Monkey achieves a score of 802, outperforming 8B-parameter state-of-the-art model InternVL2-8B. Besides, our model and training strategy are very efficient, which can be trained with only eight RTX 3090. The code is available at https://github.com/Yuliang-Liu/Monkey.
footnotetext: Y. Liu, D. Liang, and X. Bai are with Huazhong University of Science and Technology. M. Huang and L. Jin are with South China University of Technology. This work was done when M. Huang was visiting Huazhong University of Science and Technology. Corresponding authors.
1 Introduction
In recent years, the field of natural language processing (NLP) has demonstrated a significant paradigm shift, marked by a focus on the development of Large Language Models [77, 3, 64, 54] (LLMs). This shift has paved the way for the creation of multimodal large language models (MLLMs) capable of processing general vision-and-language understanding [32, 40, 2]. Researchers are actively exploring effective and efficient methods for integrating vision encoders with LLMs. Some methods, such as Flamingo [1], BLIP-2 [32], MiniGPT4 [79], and Qwen-VL [2] utilize a set of learnable queries to sample the image tokens and align the image tokens with Large Language Models. In contrast, other methods like LLaVA [41] and CogVLM [65] propose to use a linear layer to achieve this. Despite these achievements, detailed scene understanding was not achieved well by previous multimodal large language models due to the limited resolution to handle.
Figure 1: Sawtooth Effect caused by the cropping. (a) Input Image (b) Cropping strategy. (c) Overlapping Cropping Strategy. (d) Ours: Multi-scale adaptive crop strategy.
Recent efforts have attempted to tackle this issue by expanding the input resolution of the image. The cropping strategy is one of the most commonly used methods [39, 71, 35, 7, 58, 70]. There are many technical extensions to the simplest cropping strategy. For instance, Monkey [35] leverages the LoRA [22] into the vision encoder for learning detail-sensitive features from the sub-image. Although these methods have shown promising results, their performance still lags behind leading commercial models. To bridge this gap, InternVL 1.5 [7] employs a powerful vision encoder [8] to enhance visual representation and uses dynamic high resolution to scale up the resolution to 4K, significantly improving the performance.
Despite the significant progress achieved by multimodal large language models, challenges in detailed scene understanding persist due to the cropping strategy. Cropping operations on images inevitably segment objects and connected areas, impairing the MLLM’s ability to recognize small or irregularly shaped objects, particularly in the context of document understanding. This strategy will introduce two types of semantic incoherence: 1) If an object or character is divided, it may not be recognized [24]. For instance, after cropping, the nose looks very much like a monkey, as shown in Fig. 1(b); 2) If a word or sentence is segmented, the semantic damage of the segmented word will be caused. For example, the word ‘Breakdown’ may be divided into ‘Break’ and ‘down’, causing semantic damage to the segmented word [44, 76]. For simplicity, we call this issue the sawtooth effect in this paper. A very straightforward idea is to adopt an overlapping cropping strategy to solve this issue, as presented in Fig. 1(c). However, as presented in our ablation studies Sec. 4.3, the overlapping cropping strategy introduces certain hallucinations that cause performance to decrease rather than increase. Moreover, this sawtooth effect is particularly evident in lightweight MLLMs, as discussed in Sec. 4.4. Larger MLLMs with enhanced comprehension capabilities can alleviate this issue to some extent.
In this paper, we propose Mini-Monkey, a lightweight multimodal large language model designed to mitigate the sawtooth effect caused by cropping strategies. Unlike existing methods [71, 7, 36] that directly crop input images, Mini-Monkey employs a plug-and-play method termed multi-scale adaptive crop strategy (MSAC), which enables effective complementation between features from different scales, as shown in Fig. 1(d). MSAC first performs a stratified operation on a pre-set group of grids according to the aspect ratios and the resolution of these grids. It then adaptively selects multiple aspect ratios from each stratified layer, ensuring that the same text is not split across different images. Multiple images will be generated based on the aspect ratios and processed by a pre-trained vision encoder to generate multi-scale visual representations. These representations are concatenated into a sequence and fused within the LLM to interact with each other. With the MSAC, Mini-Monkey adaptively generates multi-scale representations, allowing the model to select non-segmented object features from various scales. The MSAC may introduce some additional computational overhead. Therefore, we propose a Scale Compression Mechanism (SCM) for use in situations where there are restrictions on computational overhead. SCM is a training-free and parameter-free module to reduce the computational overhead. It utilizes the well-trained attention layers from the LLM to produce the attention weight and dropout token based on the attention weight.
Experiments have demonstrated the effectiveness of Mini-Monkey: 1) Mini-Monkey achieves state-of-the-art performance among 2B-parameter MLLMs in both general multimodal understanding and document understanding tasks. Especially, Mini-Monkey outperforms the state-of-the-art 2B-parameter method by an average of 1.7% across 13 benchmarks in terms of evaluation metrics; 2) Surprisingly, we find that Mini-Monkey achieves a score of 802 on the OCRBench, outperforming the 8B-parameter state-of-the-art model InternVL2-8B.
2 Related Works
2.1 Multimodal Large Language Models
In recent years, Large Language Models (LLMs) have made significant progress [77, 3, 64, 54, zeng2022glm,llama32024]. Drawing from this advancement, many efforts have been made to integrate a vision encoder into Large Language Models for vision-language understanding. A commonly employed approach is the Linear Projector method[40, 65, 79], which maps the output of the vision encoder to the same feature space as the text features of the Large Language Models. Some methods, such as QFormer [32], Perceiver Resampler [1, 2], or Abstractor [72], introduce a set of learnable queries to facilitate this integration. Despite notable progress, previous methods face challenges in handling the detailed scene understanding due to the limited resolution. To address this issue, recent works have primarily employed the following strategies: 1) Two vision encoders, one for processing high-resolution images and one for processing low-resolution images [67, 78, 20]. 2) Directly using visual encoders that support high-resolution input [37, 48]. 3) Using a cropping strategy to segment the high-resolution images into several low-resolution images [71, 36]. While these methods have effectively enhanced resolution, they still display shortcomings in document understanding, especially when compared to top commercial models. To close the gap, InternVL1.5 [7] utilizes a large vision encoder [8] and a dynamic high-resolution strategy to train on high-quality data. Concurrently, LLama3-V [63] employed a cropping strategy to enhance resolution, releasing several models with varying parameter counts, reaching up to 400 billion. Although LLama3-V and InternVL1.5 achieves promising results on several multimodal benchmarks, the cropping strategy used in it will inevitably result in semantic incoherence: 1) If an object or character is divided, it may not be recognized; 2) If the word or sentence is segmented, the semantic damage of the segmented word will be caused. For example, the word ‘Breakdown’ may be divided into ‘Break’ and ‘down’, causing semantic damage to the segmented word [44, 76]. This will limit it in the detailed scene understanding. Although some methods [44, 24] attempt to address this issue by introducing attention modules, they introduce additional parameters and require training this module from scratch. In contrast, our method is plug-and-play, requiring no additional parameters.
2.2 Visually-situated document understanding
The visually-situated document understanding is a task that comprehends rich text information in the images, including natural images [59, 60], documents images [53, 61, 51], charts images [50, 26], tables images [55, 6], etc. document understanding models can be broadly categorized into two types based on their reliance on OCR systems for text extraction: OCR-dependent methods and OCR-free methods. OCR-dependent methods use the text extracted from the OCR system to perform related document understanding tasks. For instance, LayoutLM v3 [25] learns the multimodal representations by unified text and image masking pre-training objectives. UDOP [62] develops a unified framework to learn and generate vision, text, and layout modalities together. On the contrary, OCR-free methods perform the document understanding tasks in an end-to-end manner without OCR input. Dount [28] directly generates textual elements based on the document images without OCR input. Pix2Struct [30] further leverages structure information by generating the HTML DOM tree for masked screenshots of web pages. Additionally, some methods incorporate large language models (LLMs) for enhanced document understanding. Ureader [71] presents a shape-adaptive cropping module to crop input images based on their aspect ratios. Similarly, TextMonkey [44] introduces enhancements for cross-window relations and a token resampler for document understanding. LayoutLLM [48] develops LayoutCoT for layout-aware supervised fine-tuning. StrucTexTv3 [49] introduces a lightweight multimodal large language model designed for perceiving and comprehending text-rich images.
2.3 Lightweight Multimodal Large Language Models
Due to the substantial computational costs associated with multimodal large language models (MLLMs), recent efforts have focused on developing more efficient models for rapid development and real-world applications. LLaVA-Phi [80] and Imp [57] leverage a lightweight language model combined with a vision encoder to create a lightweight large multimodal model. In this context, several researchers are exploring efficient architectural designs. For instance, MobileVLM [9] introduces a lightweight downsample projector to minimize resource usage during training and inference. Bunny [19] offers an efficient data compression technique to reduce the volume of pretraining data required. TinyGPT-V [75] employs a multi-stage training approach tailored for lightweight multimodal models. Similarly, MiniCPM presents a scalable training strategy aimed at producing an efficient lightweight multimodal large language model. Additionally, Vary-toy [68] supports high-resolution input, while InternVL 2 [7] enhances the performance of lightweight MLLMs through a dynamic high-resolution strategy. Despite these promising advancements, the state-of-the-art method faces limitations due to a sawtooth effect caused by the cropping strategy.
Figure 2: The overall architecture of Mini-Monkey. H-Attn represents high attention weight. L-Attn represents low attention weights. The tokens with low attention weights will be filtered. The shared LLM’s Layer represents using the block layer from LLM in SCM.
3 Mini-Monkey
The overall architecture is illustrated in Fig. 2. Mini-Monkey consists of an MSAC, a vision encoder, an MLP layer, a Scale Compression Mechanism, and a Large Language Model (LLM). Initially, Mini-Monkey generates multiple images through MSAC. These images are then processed by the vision encoder and MLP layer to extract image tokens. The Scale Compression Mechanism adjusts these image tokens based on the input question and forwards them to the LLM, which subsequently generates the final answers.
3.1 Multi-Scale Adaptive Crop Strategy
Previous state-of-the-art methods [36, 7] adopt the cropping-based strategy to expand the resolution of images and segment the high-resolution images into a set of sub-images. However, this cropping strategy will lead to a sawtooth effect. To address this issue, we introduce a multi-scale adaptive crop strategy (MSAC) that achieves the synergy between images with different scales to mitigate the semantic incoherence caused by the cropping strategy. As shown in Fig. 2 (b), we generate a pre-defined set of grids. The maximum of these grids is less than $Max_{num}$.
Then, we perform a stratified operation on these grids, which are divided into three sets according to their aspect ratios. We will select one aspect ratio for each layer. Different stratified layers provide different information to the model. The detailed layer $A_{d}$ is responsible for providing detailed information. It not only limits the maximum of the sub-image but also limits the minimum of the sub-image to make the image as large as possible to make the object in the image clearer. The images generated by this layer may have a semantic inconsistency. Therefore, we utilize the adaptive layer $A_{a}$ to synergize with the detailed layer, allowing the model to select non-segmented objects from various scales. The adaptive layer will adaptively generate an aspect ratio based on the detailed layer, ensuring that the cropping lines on the detailed layer and those on the adaptive layer do not overlap. This can be formulated as follows:
$C_{Id}\cap C_{Ia}=\emptyset\,.$ (1)
where $C_{Id}$ represents the cropping lines on the detailed layer. $C_{Ia}$ represents the cropping lines on the adaptive layer. Specifically, if the aspect ratio from the detailed layer is a multiple of that from the adaptive layer, we remove it from the adaptive layer and select a new ratio. This process ensures that the detailed and adaptive layers provide distinct semantic information and visual features for the model.
We also produce a global view of the image as a low resolution using an aspect ratio 1:1{1:1}, termed global layer. After obtaining the image from three layers, these images are sent to the vision encoder to extract the features and compress the visual tokens through the scale compression mechanism. Once the visual tokens are compressed, they will be fed into the large language model to conduct the multi-scale visual representations fusion and output the results.
Multi-Scale Visual Representations Fusion. Unlike the previous work [58] that simply concatenates multi-scale features along the dimension, our approach involves the fusion of multi-scale visual representations within a large language model (LLM). Within the LLM, these multi-scale visual representations interact with each other through self-attention. By fusing features from different scales, Mini-Monkey gains an enhanced ability to comprehend visual text information.
3.2 Scale Compression Mechanism
Although the proposed MSAC significantly enhances model performance, certain scenarios may impose computational requirements. To tackle this challenge, we introduce a parameter-free token compression method called the Scale Compression Mechanism (SCM), which is used to reduce the visual tokens. A well-trained LLM can effectively select the necessary visual features based on the input question. Consequently, SCM utilizes the first and second layers of the LLM to select visual tokens without generating any additional parameters. Specifically, as shown in Fig. 2 (c), the input visual token including $V_{d}\in\mathbb{R}^{L_{1}\times C}$, $V_{a}\in\mathbb{R}^{L_{2}\times C}$, and $V_{g}\in\mathbb{R}^{L_{3}\times C}$, and the textual token $T_{t}\in\mathbb{R}^{T\times C}$ will be sent into an LLM’s Layer. $V_{d}$ represents the tokens from detailed layer. $V_{a}$ represents the tokens from adaptive layer. $V_{g}$ represents the tokens from global layer. We choose the visual token from the adaptive layer, global, and textual token to attend to the visual token from the detailed layer. The LLM’s Layer will output an attention map. The calculation of the attention can be formulated as follows:
$\mathbf{Attn_{w}}={\tt softmax}(\frac{(Cat(V_{a},V_{g},T_{t})+{\rm PE}(Cat(V_{a},V_{g},T_{t})))(V_{t}+{\rm PE}(V_{t}))^{T}}{\sqrt{D}})\,.$ (2)
Where $\rm PE$ represents the position encoding and $D$ denotes the dimension of the LLM, we proceed to obtain the attention map $\mathbf{Attn_{w}}\in\mathbb{R}^{T\times L}$. $Cat()$ represents the sequence concatenation operation. Due to the lower information density of tokens from detailed layers, we primarily focus on compressing these tokens. In contrast, visual tokens from adaptive and global layers provide the LLM with complete spatial information. After obtaining the attention map, we select the top $K$ visual features from detailed layers based on this map. Compared to FastV [5], SCM is more targeted by using tokens with high relative information density to compress tokens with low information density. The ablation study in Sec. 4.3 demonstrates the effectiveness of SCM.
4 Experiments
4.1 Implementation Details
We use a well-trained InternViT [8], MLP layers, and the InternLLM [4] from InternVL2-2B [7] as the vision encoder, connector and the LLM. Following previous work [8], we use the (448, 448) as the input resolution of InternViT. During the inference, for the detailed Layer, we set the maximum of the sub-image $Max_{num}$ as 24 and set the minimum of the sub-image as 8. For the adaptive layer, we set the maximum of the sub-image as 8 and set the minimum of the sub-image as 1.
Evaluation. Following the previous work [19, 7], we evaluate Mini-Monkey on eleven general multimodal understanding benchmarks, including MathVista testmini [47], SEED Image [31], RealWorldQA [69], AI2D test [27], POPE [34], CCBench [42], MME [16], and HallusionBench [18].
For document understanding, following the previous work [44], we employ two distinct types of metrics to assess the performance of Mini-Monkey. Initially, we leverage the standard metrics provided by the benchmarks to evaluate Mini-Monkey. For this metric, similar to  [7], we utilize benchmarks such as ChartQA [50], DocVQA [53], InfoVQA [52], TextVQA [60], and OCRBench [43]. ChartQA, DocVQA, InfoVQA, and TextVQA are widely used to assess the textual comprehension of models. OCRBench, a more recent benchmark, includes 29 datasets to provide a comprehensive evaluation of the model’s capabilities. Subsequently, we apply the accuracy metric to verify the performance. For this metric, a response from Mini-Monkey that fully captures the ground truth is considered a true positive. Further details on this metric and the used benchmarks can be referenced in [43].
Table 1: Comparison with SoTA models on 8 multimodal benchmarks. General multimodal benchmarks encompass: MME [16], RealWorldQA [69], AI2D test [27], CCBench [42], SEED Image [31], HallusionBench [18], and POPE [34]. Additionally, the math dataset includes MathVista testmini [47]. The MME results we report are the sum of the perception and cognition scores. § represents the results from the OpenCompass leaderboard [10].
General Multimodal Benchmarks Math
model #param MME RWQA AI2D CCB SEED HallB POPE MathVista
Mini-Gemini [33] 35B 2141.0 - - - - - - 43.3
LLaVA-NeXT [39] 35B 2028.0 - 74.9 49.2 75.9 34.8 89.6§ 46.5
InternVL 1.2 [8] 40B 2175.4 67.5 79.0 59.2 75.6 47.6 88.0 47.7
InternVL 1.5 [7] 26B 2187.8 66.0 80.7 69.8 76.0 49.3 88.3 53.5
DeepSeek-VL [46] 1.7B 1531.6 49.7§ 51.5§ 37.6§ 43.7§ 27.6§ 85.9§ 29.4
Mini-Gemini [33] 2.2B 1653.0 - - - - - - 29.4
Bunny-StableLM-2 [19] 2B 1602.9 - - - 58.8 - 85.9 -
MiniCPM-V-2 2.8B 1808.6 55.8§ 62.9§ 48.0§ - 36.1§ 86.3§ 38.7
InternVL 2 [7] 2B 1876.8 57.3 74.1 74.7 70.9§ 37.9 85.2§ 46.3
Mini-Monkey (ours) 2B 1881.9 57.5 74.7 75.5 71.3 38.7 86.7 47.3
Table 2: Comparison to state-of-the-art MLLMs on OCR-related Tasks. Mini-Monkey achieves the best results among the 2B-parameter MLLMs. § represents the results from the OpenCompass leaderboard [10].
Model Model Size DocVQATest ChartQATest InfoVQATest TextVQAVal OCRBench
TextMonkey [44] 9B 73.0 66.9 28.6 65.6 558
TextHawk [74] 7B 76.4 66.6 50.6
DocKylin [76] 7B 77.3 46.6 66.8
HiRes-LLaVA [24] 7B 74.7 61.5 48.0 65.4
LLaVA-UHD [70] 13B 67.7
CogAgent [20] 17B 81.6 68.4 44.5 76.1 590
UReader [71] 7B 65.4 59.3 42.2 57.6
DocOwl 1.5 [21] 8B 82.2 70.2 50.7 68.6
HRVDA [37] 7B 72.1 67.6 43.5
IXC2-4KHD [13] 8B 90.0 81.0 68.6 77.2 675
InternVL 1.5 [7] 26B 90.9 83.8 72.5 80.6 724
InternVL 2 [7] 8B 91.6 83.3 74.8 77.4 794
GLM4-V [17] 9B - - - - 786
Vary-toy [68] 1.8B 65.6 59.1 - - -
MiniCPM-V 2.0 2.8B 71.9 55.6§ - 74.1 605
InternVL 2 [7] 2B 86.9 76.2 58.9 73.4 784
Mini-Monkey (Ours) 2B 87.4 76.5 60.1 75.7 802
Table 3: Quantitative accuracy (%) comparison of our model with existing multimodal large language models (MLLMs) on several benchmarks. Following TextMonkey [44], we use the accuracy metrics to evaluate our method.
Method Scene Text-Centric VQA Document-Oriented VQA KIE
STVQA TextVQA DocVQA InfoVQA ChartQA FUNSD SROIE POIE
BLIP2-OPT-6.7B [32] 20.9 23.5 3.2 11.3 3.4 0.2 0.1 0.3
mPLUG-Owl [72] 30.5 34.0 7.4 20.0 7.9 0.5 1.7 2.5
InstructBLIP [11] 27.4 29.1 4.5 16.4 5.3 0.2 0.6 1.0
LLaVAR [78] 39.2 41.8 12.3 16.5 12.2 0.5 5.2 5.9
BLIVA [23] 32.1 33.3 5.8 23.6 8.7 0.2 0.7 2.1
mPLUG-Owl2-8 [73] 49.8 53.9 17.9 18.9 19.4 1.4 3.2 9.9
LLaVA1.5-7B [38] 38.1 38.7 8.5 14.7 9.3 0.2 1.7 2.5
TGDoc [66] 36.3 46.2 9.0 12.8 12.7 1.4 3.0 22.2
UniDoc [15] 35.2 46.2 7.7 14.7 10.9 1.0 2.9 5.1
DocPedia [14] 45.5 60.2 47.1 15.2 46.9 29.9 21.4 39.9
Monkey-8B [36] 54.7 64.3 50.1 25.8 54.0 24.1 41.9 19.9
InternVL-8B [8] 62.2 59.8 28.7 23.6 45.6 6.5 26.4 25.9
InternLM-XComposer2-7B [12] 59.6 62.2 39.7 28.6 51.6 15.3 34.2 49.3
TextMonkey-9B [44] 61.8 65.9 64.3 28.2 58.2 32.3 47.0 27.9
Mini-Monkey-2B (Ours) 66.5 68.4 78.1 49.6 57.9 42.9 70.3 69.9
Table 4: Ablation study of multi-scale adaptive crop strategy. We compare our method with the existing crop strategy and the overlay cropping strategy.
Model Resolution Strategy TextVQA OCRBench MME HallB POPE
Baseline Dynamic High-Resolution strategy [7] 73.4 784 1876.8 37.9 85.2
Baseline Fixed Size High-Resolution strategy [36] 74.2 772 1824.5 37.6 85.0
Baseline Overlapping Cropping Strategy 70.6 758 1874.1 36.8 83.5
Baseline Multi-Scale Strategy [58] 74.8 776 1846.8 38.1 85.3
Mini-Monkey (Ours) Multi-Scale Adaptive Crop Strategy 75.7 802 1881.9 38.7 86.7
4.2 Comparison to the State of the Art
General Multimodal Understanding. We evaluate Mini-Monkey on general multimodal understanding following  [19, 7]. The results are shown in 1. Mini-Monkey surpasses other 2B-parameter models on 11 benchmarks. On the MathVista and POIE, Mini-Monkey outperforms the previous state-of-the-art method InternVL2-2B by 1% and 1.5%, respectively. On the HallusionBench, Mini-Monkey outperforms MiniCPM-V-2 by 2.6%. These results showcase the ability of Mini-Monkey to handle general multimodal understanding and reasoning tasks.
document understanding. For the first type of metric, the results are presented in Tab. 2. We use a relaxed accuracy measure for ChartQA, ANLS for DocVQA and InfoVQA, and the VQA score for TextVQA. The results indicate that Mini-Monkey achieves state-of-the-art performance among 2B-parameter multimodal large language models. Compared to InternVL2-2B, our method outperforms it by 2.3%, 1.8%, and 1.2% for TextVQA, InfoVQA, and OCRBench, respectively. Due to the small original resolution of ChartVQA, it is less impacted by cropping operations, resulting in a minor improvement from our method. Notably, in the OCRBench, Mini-Monkey even surpasses the 8B-parameter Large Multimodal Model InternVL2-8B and the 9B-parameter Large Multimodal Model GLM4-V by 0.8% and 1.6%, respectively. These results demonstrate the advantages of a multi-scale adaptive crop strategy in enhancing document understanding.
For the accuracy metric, the results are shown in Tab. 3. Mini-Monkey shows an average performance improvement of 14.8% compared to TextMonkey-9B [44], demonstrating the effectiveness of our method. Mini-Monkey also outperforms the state-of-the-art methods on multiple text-related benchmarks. Specifically, Mini-Monkey achieves 49.2% on FUNSD, 70.3% on SROIE, and 69.9% on POIE, outperforming the previous state-of-the-art method by 10.6%, 23.3%, and 42.0%, respectively. These results further indicate the great potential of Mini-Monkey for downstream task applications, such as visual key information extraction.
[…]
Model MSAC TextVQA OCRBench MME HallB POPE
[…]
MiniCPM-V-2 $\checkmark$ 76.0 (+1.9) 627 (+22) 1819.5 (+10.9) 36.5 (+0.4) 87.1 (+0.8)
InternVL 2 $\times$ 73.4 784 1876.8 37.9 85.2
InternVL 2 $\checkmark$ 75.7 (+2.3) 802 (+18) 1881.9 (+5.1) 38.7 (+0.8) 86.7 (+1.5)
Table 6: Ablation study of the scale compression mechanism. We used different compression ratios to compare with FastV [5]. (0.5) represents 50% compression and (0.9) represents 90% compression.
Model Resolution Strategy TextVQA OCRBench MME HallB POPE
Mini-Monkey Pooling (0.5) 47.6 256 1765.2 31.5 84.5
Mini-Monkey Random (0.5) 63.5 503 1805.5 36.2 85.9
Mini-Monkey FastV [5] (0.5) 73.4 781 1848.0 38.3 83.9
Mini-Monkey FastV [5] (0.9) 73.9 792 1866.1 37.5 85.8
Mini-Monkey SCM (0.5) 74.7 794 1886.0 38.7 86.1
Mini-Monkey SCM (0.9) 75.2 801 1884.7 38.6 86.2
4.3 Ablation Study
In this section, we perform ablation studies on both general multimodal understanding and document understanding benchmarks to validate the effectiveness of our method. We adopt the TextVQA [60], OCRBench [43], HallusionBench [18], MME [16], and RealWorldQA [69] to conduct ablation studies.
Multi-Scale Adaptive Crop Strategy. We conducted ablation studies to investigate the effectiveness of the proposed multi-scale adaptive crop strategy. We compared our method with several alternatives: The dynamic high-resolution strategy [7], which maintains aspect ratios to increase resolution. The fixed-size high-resolution strategy [36], which uses a fixed size to increase resolution. The overlapping cropping strategy, which uses a high-resolution approach but crops with overlay. The multi-scale strategy [58], which introduce a fixed size multi-scale strategy to the MLLM. As presented in Tab. 4, the proposed multi-scale adaptive crop strategy achieved the best results. Our method outperforms the fixed-size multi-scale strategy [58] by 3.3%. Notably, the over-overlay cropping strategy, instead of improving the model’s performance, actually degraded it.
The proposed multi-scale adaptive crop (MSAC) strategy can be seamlessly integrated into crop-based methods. To demonstrate its effectiveness, we incorporated MSAC into various structures of MLLM, such as MiniCPM-V-2 and InternVL 2. As shown in Tab. 5, MSAC consistently enhances the performance across different MLLM structures, thereby validating the effectiveness of our approach.
Scale Compression Mechanism. We compared the proposed Scale Compression Mechanism with the related work FastV [5]. For different methods, we compress the number of visual tokens by 50%. For our method and FastV, we further conduct an experiment with 90% compression. Following FastV’s paper, we set the K in FastV as 2. FastV is a plug-and-play method. Therefore, we conducted this experiment without training the model. As illustrated in Tab. 6, when using 50% compression and 90% compression, our method outperformed FastV by 21.5% and 4.4%, respectively, demonstrating its effectiveness. Both our method and FastV are parameter-free. FastV compresses input tokens, including both visual and textual tokens, within Transformer blocks. In contrast, our method is more targeted by using tokens with high relative information density to compress tokens with low information density.
Figure 3: Qualitative results of Mini-Monkey. (a) Input Image and Ground Truth. (b) The results of using overlapping cropping strategy. OSC represents overlapping cropping strategy. (c) The results of InternVL2-2B and InternVL2-26B. (d) The results of Mini-Monkey.
4.4 Qualitative Results
In this section, we provide some qualitative results to demonstrate the effectiveness of our method. We first verify that the sawtooth effect is particularly evident in lightweight MLLMs, adopting InternVL2-2B and InternVL2-26B. As shown in Fig. 3(c), InternVL2-26B can answer the questions correctly. However, due to the word ‘classrooms’ and ‘school’ being cropped, InternVL2-2B gives a wrong answer that addresses the text in the bottom left corner of the original image. While Mini-Monkey can overcome this sawtooth effect and provide the correct answer, as presented in Fig. 3(d). Comparing Fig. 3(b) and Fig. 3(d), we can see that the overlapping cropping strategy introduces some hallucinations and cannot answer questions accurately based on the image, whereas our methods can effectively address the sawtooth effect.
5 Discussion
There are some other methods to address this issue. One approach is to use a vision encoder that inherently supports high resolution, such as the Swin-Transformer [45] or SAM [29]. However, the cropping strategy remains the most commonly employed method. This preference stems from the ability to leverage the pre-trained, robust vision encoder CLIP [56]. Why high-resolution encoders are not always used to tackle this problem directly? The reason lies in the resource efficiency of CLIP pre-training. Typically, due to the low-resolution input, CLIP requires fewer resources compared to high-resolution visual encoders. Consequently, CLIP is often chosen as the visual encoder in multimodal large language model (MLLM) systems, with the cropping strategy being used to enhance the input resolution.
6 Conclusion
In this study, we introduced Mini-Monkey, a lightweight multimodal large language model (MLLM) designed to address the limitations of existing cropping strategies used to enhance MLLMs’ ability to process high-resolution images. Traditional cropping methods often segment objects and connected areas, which limits the recognition of small or irregularly shaped objects and text, a problem particularly pronounced in lightweight MLLMs. To mitigate this, Mini-Monkey employs a multi-scale adaptive crop strategy (MSAC), generating multi-scale representations that allow for the selection of non-segmented objects across different scales. The proposed MSAC can be consistently enhanced across various MLLM architectures. Additionally, we developed a Scale Compression Mechanism (SCM) to reduce the computational overhead of MSAC by compressing image tokens. Our experimental results demonstrate that Mini-Monkey not only achieves leading performance on a variety of general multimodal model understanding tasks but also shows consistent improvements in document understanding tasks. Notably, on the OCRBench benchmark, Mini-Monkey scored 802, surpassing larger 8B-parameter state-of-the-art models like InternVL2-8B. Furthermore, our model and training strategy are exceptionally efficient, requiring only eight RTX 3090 GPUs for training. These results indicate the potential of Mini-Monkey as a powerful and efficient solution for advancing multimodal large language models capabilities in high-resolution image processing.